OpenStudy (anonymous):

Can someone factor this please? 2 sin^2 x = 2 + cos x interval [0, pi]

OpenStudy (gorv):

sin^2x=1-cos^2x

OpenStudy (anonymous):

factor as in simplify or factor out sin cos?

OpenStudy (anonymous):

2sin^2 x=2 +cos x 1.)Factor, Use the Trig identity 1-sin^2x=1-cos^2 x 2(1-sin^2x)=2+ cos x 2.) 2(1-cos^2x)=2+cosx 3.) 2-2cos^2x=2+cosx 4.) 2-2cos^2x-2cosx=0 5.) -2cos^2x-cosx=0 6.) 2cos^2x + cosx=0 7.) 2cos^2x+1=0 8.) cosx=-1/2 9.) x= pi/2, 2pi/3, 4pi/3, 3pi/2

OpenStudy (anonymous):

Let me know if you understand

OpenStudy (anonymous):

@n648c788 yes, completely understood. thanks so much.